# Model Compression in Practice: Lessons Learned from Practitioners Creating On-device Machine Learning Experiences

### v1

@@ 5.4.1. Budgeting by Technical Feasibility
“At the beginning of the design, [we do a] ‘back of envelope calculation’ of where things need to be as honest as possible to what’s realistic. At the beginning, you’re more focused on accuracy. Over time there is refinement.” — E7
A common refinement practice is outlined in Figure 1. First, ML engineers will work on a “feasibility” model simply to see if the ML task “is possible at all” [P3]. Once the feasibility model works at sufficient accuracy (Figure 1A), engineers work on an architecture search to find a smaller, more efficient model that achieves the accuracy goal (Figure 1B) [P3, E7]. Next, the budget is refined based on the potential to shrink the model using compression: “given [an] accuracy goal, what is the biggest percentage reduction possible?” [E11]. Some teams do this reduction estimation on their own, while others bring in ML compression experts [T5]. Since the compressed-model budget (Figure 1C) is often estimated before engineers embark on model compression, the budget remains open to refinement, subject to product design constraints. Ultimately if an ML-powered feature requires a strict model budget, the model will not ship until engineers have found a way to reach those thresholds [P3].
5.4.2. Budgeting by User’s Experience of Model
@@
[Uncaptioned image] [Uncaptioned image] One-Time, Real-Time, or All the Time?
Latency budget is almost always a per-feature UX decision dependent on how a user perceives the model output and how often the model needs to run. For example, when tagging people in a new photo, the model needs to work fast enough that users do not notice a delay. Models on-device remove the wait on network response times, which make these latency budgets easier to achieve [P2].
Models that only need to run once generally have a looser resource budget than continuous models. For models running in real-time, only a few milliseconds per evaluation may be the maximum budget [P1]. This is most crucial in real-time scenarios where the user can perceive the model working, which often occurs in live video applications [E13]. A user will notice if their video feed appears choppy or a background filter appears delayed. Latency budget needs to be as low as possible. Equally, if not more, strict latency budgets are given to models that need to run all the time, i.e., “always-on”. Always-on models are continuously using the device’s resources, so they are required to be as quick and lightweight as possible. Some always-on models require a timely response to the user, for example, an always-on model that listens for a user to trigger a voice-assistant.
@@
There are some applications where a ML model is the main enabling technology. For example, if the user explicitly asks a voice assistant for help or uses an app to translate languages. If a user explicitly presses a button to trigger a model, the user is in control of when the model runs (and does not run). Product designers will allow models a bigger budget for power and compute resources when the user explicitly opts-in [E11, E7, P2]. It is also a common scenario to have models running support functions in the background. Similar to any other background task, these models require a strict budget to stay unobtrusive:
@@
Many devices today will download and install their routine updates at nighttime, or whenever the device’s user sleeps and has their device connected to power. Updates are intensive computational tasks that can be disruptive to normal device function, thus it is conventional to wait until a user is not actively using their device. By the same logic, scheduling a model to run overnight is a good option if the model’s output is not something needed immediately (e.g., updating tags of a photo library). If the user’s device is plugged in, this also alleviates the concern of a model draining the battery. Disruption to the user is minimized, so overnight models are also free to take a longer time with a bigger resource budget [E13, P3].
@@ 5.4.3. Top-down Budget by Application & Device
The difficulty of hitting budget can vary enormously depending on the model. Thus negotiation over budget allocations can continue and evolve until late in development [P1].
@@
Figure 2. Models needs to share space with the rest of an application. Model budget typically reflects how “valuable” a model is compared to all other features of the application.
@@
A model’s power budget is a measure of how much it is permitted to drain a device’s battery life. Power consumption of a model can be tricky to estimate, and should be measured empirically on the target device for the most reliable estimate [E6, T4, P9]. Once again, real-time or always-on models require much stricter power budgets, because they continuously draw power. Instead of measuring power at one moment of execution, these continuous models can be thought of in terms of the total percentage or total minutes of battery they consume in a day [E6, P3, E7].
[Uncaptioned image] Thermal Throttling
Closely related to power consumption is device temperature. In the words of E7: “ship high-quality ML models that don’t melt the user’s device.” By melting, E7 refers to both a device that is uncomfortably warm to the touch, and a device that has been forced into thermal mitigation. Thermal mitigation, or thermal throttling, is when a device slows down processing to protect itself from overheating. Hot and slow devices is a terrible user experience that must be avoided. Thus, memory and power budgets must be set according to thermal measurements [E6, E7, E4, E1, E2]: “Heat throttling is everyone’s budget!” [E4].
@@ 5.4.5. Gating & Variable-Budget Options
“So people think about model compression as static resources, but the actual resources on device are dynamic based on what’s running. So it’s useful to be able to on-the-fly change how much resources your model is uses.” — E10
For any model, there will be some size and efficiency budget at which the model can be shrunk without any noticeable difference to the user. Below that threshold, accuracy degradation may degrade UX. A budget must balance the best model accuracy with the lowest possible resource footprint.
Strategy #3: On-device ML does not need to be all-or-nothing. If current compression techniques cannot fit your ML feature on-device, try breaking the feature down into subtasks. Delegate smaller subtask models to live on-device and delegate larger ML workloads to a server. Prioritize on-device ML for feature subtasks that will help preserve your user’s privacy and keep critical functionality responsive in the absence of a network connection.
5.5. Considerations for Investing in Compression
Participants strongly emphasized that model compression is an investment. Compression techniques can take intensive engineering effort to apply—and do not always work. There are many circumstances that can cause one compression technique to generate enormous savings in one model, and completely fail to produce any savings in a different model, potentially after weeks of wasted engineering effort. While this work is not an exact science, we detail general tips and guidance from practitioners. Note that most of these techniques assume the model to compress is a neural network.
@@ 5.5.1. Maximize Architecture Savings First
As shown in Figure 1, practitioners often start optimizing a model by finding a smaller architecture first. “If you’re running a common ResNet, look at strides, layer widths, number of layers” [T5]. Participants also recommended neural architectures designed for device efficiency [E13, P5, E3, E11, T6], such as MobileNet (Howard et al., 2017). Some applications do not need a heavy neural network; instead it might be possible, and even beneficial, to convert into a simpler decision tree or SVM [P3]. Architecture savings are seen as a more reliable first approach before attempting compression [E13, E1]:
@@
Due to the high training cost of some models, participants discussed manual architecture selection (as opposed to auto-ML solutions that automate various stages of ML development) based on their expertise: “It’s really trial and error by hand monitoring the hardware usage [and] identifying bottlenecks,” [E2]
@@ 5.5.2. Check the Architecture Against the Hardware
A critical caveat to architecture optimizations is that it is easy to be fooled. A common mistake is to trust that academic efficiency results will be reproducible: “Just because you read a paper, don’t expect what is “efficient” in a paper to exactly work in practice,” [E7]. The issue is that at the lowest level of computation, different hardware optimizes for different operations. Efficiency measures reported in papers are only reliable for the specific hardware set they were run on. Even for specialized accelerator hardware designed for ML models and neural networks, the hardware will likely only support certain types of operations. Since hardware cannot be altered at the same pace as software, the latest neural layer architectures from papers may not be possible to run efficiently on-device. Practitioners emphasized the need to adapt their neural architecture to better fit the hardware [T5, E7, P9].
@@ 5.5.3. Test Against a Range of Hardware
In conventional machine learning work, and more generally much of modern software engineering, it is not required to know the details of the hardware that code will run on. In efficient machine learning, hardware may be one of the greatest challenges when working in on-device machine learning:
@@
To further complicate matters, many times practitioners are not considering a single specific hardware implementation, but rather a class of hardware. Classes of hardware include platforms such as mobile phones, smartwatches, tablets, computers, and others. Consider a practitioner building a model for a smartphone. There are many types of smartphones, and versions of existing smartphones each with their own memory and compute details. Hardware also changes over time, and practitioners may need to consider older versions of hardware that have already been shipped. In the same way that front-end developers build responsive UIs and applications for multiple screensizes, so too do ML practitioners now need to optimize and test multiple hardware implementations specific to a set of devices [E6].
@@
Practitioners we spoke with had a sense of “accuracy v. effort” [E6] or “risk v. reward” [E13] for some of the most common compression approaches. As a general heuristic, compression techniques that need little-to-no retraining will be cheaper to apply—but at the cost of steeper model accuracy degradation. Since some of these models take days to train, involving complex compression logic in the training process is costly. Practitioners must consider cost in money, time, and effort that implementing a specific compression technique could take.
@@
When model accuracy goes down after quantization or pruning, fine tuning can help recover accuracy (Gholami et al., 2021; Hoefler et al., 2021). Fine tuning a model costs training time, but is necessary in many situations to recover accuracy. Pruning techniques almost always require fine-tuning on the pruned model (Hoefler et al., 2021). 6/30 participants discussed successfully utilizing this approach [P8, T3, T1, E8, E13, E10].
@@
Though it introduces additional complexity to model training, compression can be approached as yet another optimization that a model needs to learn during training. 11/30 participants discussed having experience with quantization learned during training [E6, E9, T5, E3, E5, E7, E10, E1, E12, T2, P10]. This approach is generally considered best-in-class and often required when designing always-on ML experiences that require low latency, e.g., real-time computational photography models. Practitioners advised training-aware compression when the model needs to be compressed significantly to meet budget while keeping high accuracy:
@@
Practitioners had less experience with learned sparsity techniques, calling learned sparsity “high risk high reward” [E13] because it is much harder to control than post-training sparsity techniques. Although training-aware compression is considered the best form of compression for the greatest model shrinkage with the least accuracy degradation (Gholami et al., 2021), a major drawback is that is must be planned for with initial model training: “Not starting early with compression is a dead end,” [E3].
@@
“[It] gives you a sense of how much quantization is going to help at all before you go through expensive training step.” — T5
If post-training techniques like quantization produce results that are too far from budget goals, teams can then pivot to training-aware approaches.
@@
5.5.5. Determine if a Model is Memory Bound or Compute Bound
If memory-bound, architecture changes, sparsity techniques, or Palettization techniques can help reduce size on-device. If the data size is causing the memory issue, reducing data resolution for model processing can help [E8], e.g., reducing video data from 1080p to 720p. On-disk storage is considered the easiest memory savings to get from model compression [E13].
If compute-bound, techniques like quantization can help speed up and simplify complex matrix math. Quantization can also save memory since smaller numbers help keep computation in cache memory [E12]. Cache locality is a common latency bottleneck to check [E6, E12], that also helps with power problems:
@@ 5.6. Evaluating Compressed Models: Efficiency, Accuracy, and Artifacts
The goal of model compression is to reduce a model’s size while preserving its accuracy, or what some refer to as “acceptable accuracy degradation” [E8]. In general, past a certain point, shrinking a model will likely degrade its accuracy. In literature, there are examples showing that compression can actually improve accuracy, by acting as a model regularizer and forcing the model to generalize better (Kusupati et al., 2020; Touvron et al., 2021). This phenomenon is sometimes referred to as Occam’s Hill (Hoefler et al., 2021): as light compression is applied, there can be an slight accuracy bump as the model is forced to generalize; however, as one increases the compression strength the model becomes too general to properly work and accuracy quickly drops. While these results have been observed on academic benchmarks, practitioners said that their models tend to generalize quite well already, so most often they see little improvement from the regularizer effect of compression [E8, E1]. In addition, many practitioners start with very large models they need to shrink models past the point of observing accuracy boosts from compression. Thus, we focus this discussion on the much more common scenario where practitioners must wrestle with accuracy degradation.
@@ 5.6.1. The Trade-off Curve Visualization
To empirically compare multiple compressed models and select the one that satisfies budget requirements, practitioners typically plot a model behavioral metric (e.g., accuracy) on the y-axis against any performance metric (e.g., model size, model latency, or power consumption) on the x-axis (Vasu et al., 2023b, a). This is illustrated in Figure 3. Each dot represents a model architecture and compression pair. These charts are called multiple names throughout different teams, including the trade-off curve, the an accuracy Pareto curve, or the tuning curve.
P3 described a scenario when optimizing over 6,000 small models. In their application, the ML model was already on the order of kilobytes; therefore, they could retrain many new models quickly. Comparing thousands of models in parallel, they plotted the F1-score on the y-axis against the model size on the x-axis. This arrangement formed the usual curve. Teams then filtered out models that do not satisfy their budgets, and ultimately selected a single models in the “knee” of the curve that has the best balance between what metrics they care about (e.g., Pareto optimal).
@@
Two scatter plots where each dot represents a compressed ML model. The y-axis encodes model accuracy from 0%–100%. The x-axes of each chart encode latency (in ms) and power consumption (in mw) respectively. Each chart shows a slight curve from the bottom-left to the top-right, showing a trend that as a model gets more accurate it tends to have a higher latency and larger power consumption. The models that are chosen to be the best are the in knee of the curve and balance accuracy, latency, and power.

### v2

@@ 5.4.1. Budgeting by Technical Feasibility
“At the beginning, [we do a] ‘back of envelope calculation’ where things need to be as honest as possible to what’s realistic. At the beginning, you’re more focused on accuracy. Over time there is refinement.” — E7
A common refinement practice is outlined in Figure 1. First, ML engineers will work on a “feasibility” model simply to see if the ML task “is possible at all” [P3]. Once the feasibility model works at sufficient accuracy (Figure 1A), engineers work on an architecture search to find a smaller, more efficient model that achieves the accuracy goal (Figure 1B) [P3, E7]. Next, the budget is refined based on the potential to shrink the model using compression: “given [an] accuracy goal, what is the biggest percentage reduction possible?” [E11]. Some teams do this reduction estimation on their own, while others bring in ML compression experts [T5]. Since the compressed-model budget (Figure 1C) is often estimated before engineers embark on model compression, the budget remains open to refinement, subject to product design constraints. Ultimately, if an ML-powered feature requires a strict model budget, the model will not ship until engineers have found a way to reach those thresholds [P3].
5.4.2. Budgeting by User’s Experience of a Model
@@
[Uncaptioned image] One-Time, Real-Time, or All the Time?
Latency budget is typically a per-feature UX decision dependent upon how a user perceives the model output and how often the model needs to run. For example, in a photography app when tagging people in a new photo, a model needs to work fast enough so that users do not notice a delay. Models on-device remove the wait on network response times, which make these latency budgets easier to achieve [P2].
Models that only need to run once generally have a looser resource budget than continuous models. For models running in real-time, only a few milliseconds per inference may be the maximum budget [P1]. This is most crucial in real-time scenarios where the user can perceive the model working, such as in live video applications [E13]. A user will notice if their video feed appears choppy or a background filter appears delayed. For this reason, latency budgets need to be as low as possible. Similarly, strict latency budgets are given to models that need to run all the time, i.e., “always-on”. Always-on models are continuously using the device’s resources, so they are required to be as quick and lightweight as possible. Some always-on models require a timely response to the user, for example, an always-on model that listens for a user to trigger a voice-assistant.
@@
There are some applications where ML is the primary enabling technology, such as voice assistants or language translation apps. If a user explicitly takes action to trigger a model, the user is in control of when the model runs (or does not run). Product designers will allow models a larger budget for power and compute resources when a user explicitly opts-in [E11, E7, P2]. It is also common to have supporting models running in the background. Similar to any other background task, these models require a strict budget to stay unobtrusive:
@@
Many devices today will download and install routine updates at night, or whenever the device’s user sleeps and has their device connected to power. Updates are intensive computational tasks that can be disruptive to normal device function, thus it is conventional to wait until a user is not actively using their device to initiate. By the same logic, scheduling a model to run overnight is a good option if the model’s output is not something needed immediately. If a user’s device is plugged in, this also alleviates the concern of a model draining battery. Disruption to the user is minimized, so models running overnight are also free to take a longer time with a bigger resource budget [E13, P3].
@@ 5.4.3. Top-down Budget by Application & Device
The difficulty of hitting budget can vary enormously depending on the model. Thus, negotiation over budget allocations can continue and evolve until late in development [P1].
@@
Figure 2. Models needs to share space with the rest of an application. Model budget typically reflects how “valuable” a model is compared to all other features of the application. The above hypothetical example shows how much budget a model (blue) has in the overall system.
@@
A model’s power budget is a measure of how much battery it is permitted to use. Power consumption of a model can be tricky to estimate, and should be measured empirically on the target device for the most reliable estimate [E6, T4, P9]. Once again, real-time or always-on models require much stricter power budgets, because they continuously draw power. Instead of measuring power at one moment of execution, these continuous models can be thought of in terms of the total percentage or total minutes of battery they consume in a day [E6, P3, E7].
[Uncaptioned image] Heat Budget
Closely related to power consumption is device temperature. In the words of E7: “ship high-quality ML models that don’t melt the user’s device.” By melting, E7 refers to both a device that is uncomfortably warm to the touch, and a device that has been forced into thermal mitigation. Thermal mitigation, or thermal throttling, is when a device slows down processing to protect itself from overheating. Hot and slow devices makes for a terrible user experience that should be avoided. Thus, memory and power budgets must be set according to thermal measurements [E6, E7, E4, E1, E2]: “Heat throttling is everyone’s budget!” [E4].
@@ 5.4.5. Gating & Variable-Budget Options
“So people think about model compression as static resources, but the actual resources on device are dynamic based on what’s running. So it’s useful to on-the-fly change how much resources your model uses.” — E10
For any model, there will be some size and efficiency budget at which the model can be shrunk without any noticeable difference to the user. Below that threshold, accuracy degradation may negatively impact UX. A budget must balance the best model accuracy with the lowest possible resource footprint.
Strategy #3: On-device ML does not need to be all-or-nothing. If current compression techniques cannot fit your ML feature on-device, try breaking the feature down into subtasks. Delegate smaller subtask models to live on-device and delegate larger ML workloads to a server if appropriate. Prioritize on-device ML for feature subtasks that will help preserve your user’s privacy and keep critical functionality responsive in the absence of a network connection.
5.5. Applying Compression during Development
Participants strongly emphasized that applying model compression is an investment. Compression techniques can take intensive engineering effort to apply—and may fail. There are many circumstances that can cause one compression technique to generate enormous savings in one model, and completely fail to produce any savings in a different model, potentially after weeks of wasted engineering effort. While this work is not an exact science, we detail general tips and guidance from practitioners. Note that most of these techniques assume the model to compress is a neural network.
@@ 5.5.1. Maximize Architecture Savings First
As shown in Figure 1, practitioners often start optimizing a model by finding a smaller architecture first. “If you’re running a common ResNet, look at strides, layer widths, and the number of layers” [T5]. Participants also recommended neural architectures designed specifically for device efficiency [E13, P5, E3, E11, T6], such as MobileNet (Howard et al., 2017). Some applications do not need a heavy neural network; instead it might be possible, and even beneficial, to convert into a simpler decision tree or SVM [P3]. Architecture savings are seen as a more reliable first approach before attempting compression [E13, E1]:
@@
Due to the high training cost of some models, participants discussed manual architecture selection (as opposed to auto-ML solutions that automate various stages of ML development) based on their expertise: “It’s really trial and error by hand monitoring the hardware usage [and] identifying bottlenecks,” [E2].
@@ 5.5.2. Check the Architecture Against the Hardware
A critical caveat to architecture optimization is that it is easy to be fooled. A common mistake is to trust in the reproducibility of academic results: “Don’t expect what is “efficient” in a paper to exactly work in practice,” [E7]. The issue is that at the lowest level of computation, different hardware optimizes for different operations. Reported efficiency measures are only reliable for the specific hardware set they were run on. Even for specialized accelerator hardware designed for ML, the hardware will likely only support certain types of operations. Since hardware cannot be altered at the same pace as software, the latest neural layer architectures from papers may not be possible to run efficiently on-device. Practitioners emphasized the need to adapt their neural architecture to better fit hardware [T5, E7, P9].
@@ 5.5.3. Test Against a Range of Hardware
In conventional machine learning work, and more generally much of modern software engineering, it is not required to know the details of the hardware that code will run on. In efficient machine learning, hardware may be one of the greatest challenges when working in on-device ML:
@@
To further complicate matters, often practitioners are not considering a single specific hardware implementation, but rather a class of hardware. Classes of hardware include platforms such as mobile phones, smartwatches, tablets, computers, and others. Consider a practitioner building a model for a smartphone. There are many types of smartphones, and versions of existing smartphones each with their own memory and compute details. Hardware also changes over time, and practitioners may need to consider older versions of hardware that have already been shipped. In the same way that front-end developers build responsive UIs and applications for multiple screensizes, so too do ML practitioners now need to optimize and test multiple hardware implementations specific to a set of devices [E6].
@@
5.5.4. Determine if a Model is Memory Bound or Compute Bound
If memory-bound, architecture changes, sparsity techniques, or palettization techniques can help reduce model size. On-disk size is considered the easiest memory savings to get from model compression [E13]. If data transfer size is causing the memory issue, reducing data resolution for model processing can help, e.g., reducing video data from 1080p to 720p [E8]. If compute-bound, techniques like quantization can help speed up and simplify complex matrix math. Quantization can also save memory since smaller numbers help keep computation in cache memory [E12]. Cache locality is a common latency bottleneck to check [E6, E12], that also helps with power problems:
@@ 5.5.4. Determine if a Model is Memory Bound or Compute Bound
Practitioners we spoke with had a sense of “accuracy v. effort” [E6] or “risk v. reward” [E13] for some of the most common compression approaches. As a general heuristic, compression techniques that need little-to-no retraining will be cheaper to apply—but at the cost of steeper model accuracy degradation. Since some of these models take days, or more, to train, involving complex compression logic in the training process is costly. Practitioners must consider cost in money, time, and effort that implementing a specific compression technique could take.
@@
When model accuracy goes down after quantization or pruning, fine tuning can help recover the loss (Gholami et al., 2021; Hoefler et al., 2021). Fine tuning a model costs training time, but is necessary in many situations to recover accuracy. Pruning techniques almost always require fine-tuning on the pruned model (Hoefler et al., 2021). 6/30 participants discussed successfully utilizing this approach [P8, T3, T1, E8, E13, E10].
@@
Though it introduces additional complexity to model training, compression can be approached as yet another optimization that a model needs to learn during training. 11/30 participants discussed having experience with quantization learned during training [E6, E9, T5, E3, E5, E7, E10, E1, E12, T2, P10]. This approach is generally considered best-in-class and often required when designing always-on ML experiences that require low latency, e.g., real-time computational photography models. Practitioners recommended training-aware compression when a model needs to be compressed significantly to meet budget while keeping high accuracy:
@@
Some practitioners had less experience with learned sparsity techniques and called them “high risk, high reward” [E13] because they are much harder to control compared to post-training sparsity techniques. Although training-aware compression is considered the best form of optimization (Gholami et al., 2021), a major drawback is that is must be included in initial model training: “Not starting early with compression is a dead end,” [E3].
@@
“[It] gives you a sense of how much quantization is going to help at all before you go through expensive training process.” — T5
If post-training techniques produce results that are too far from budget goals, teams can then pivot to training-aware approaches.
@@ 5.6. Evaluating Compressed Models: Efficiency, Accuracy, and Artifacts
The goal of model compression is to reduce a model’s size while preserving its accuracy, or what some refer to as “acceptable accuracy degradation” [E8]. In general, past a certain point, shrinking a model will likely degrade its accuracy. Conversely, in literature there are examples showing that compression can actually improve accuracy, by acting as a model regularizer and forcing the model to generalize better (Kusupati et al., 2020; Touvron et al., 2021). This phenomenon is sometimes referred to as Occam’s Hill (Hoefler et al., 2021): as light compression is applied, there can be an slight accuracy bump as the model is forced to generalize; however, as one increases the compression strength the model becomes too general to properly work and accuracy quickly drops. While these results have been observed on academic benchmarks, practitioners said that their models tend to generalize quite well already, so most often they see little improvement from the regularizer effect of compression [E8, E1]. Thus, we focus this discussion on the much more common scenario where practitioners must wrestle with accuracy degradation.
@@ 5.6.1. The Trade-off Curve Visualization
To empirically compare multiple compressed models and select the one that satisfies budget requirements, practitioners typically plot a model behavioral metric (e.g., accuracy) on the y-axis against any performance metric (e.g., model size, model latency, or power consumption) on the x-axis (Vasu et al., 2023b, a). This is illustrated in Figure 3. Each dot represents a model architecture and compression pair. These charts are called multiple names throughout different teams, including the trade-off curve, the accuracy Pareto curve, or the tuning curve.
P3 described a scenario when optimizing over 6,000 small models. In their application, the ML model was already on the order of kilobytes; therefore, they could retrain many new models quickly. Comparing thousands of models in parallel, they plotted the F1-score on the y-axis against the model size on the x-axis. This arrangement formed the usual curve. Teams then filtered out models that do not satisfy their budgets, and ultimately selected a single model in the “knee” of the curve that has the best balance between what metrics they care about (e.g., Pareto optimal).
Two scatter plots where each dot represents a compressed ML model. The y-axis encodes model accuracy from 0%--100%. The x-axes of each chart encode latency (in ms) and power consumption (in mw) respectively. Each chart shows a slight curve from the bottom-left to the top-right, showing a trend that as a model gets more accurate it tends to have a higher latency and larger power consumption. The models that are chosen to be the best are the in knee of the curve and balance accuracy, latency, and power.